# Exploiting synergies between JWST and cosmic 21-cm observations to uncover star formation in the early Universe

### v1

@@ Abstract
In the current era of JWST, we continue to uncover a wealth of information about the Universe deep into the Epoch of Reionization. In this work, we run a suite of simulations using the code 21cmSPACE  to explore the astrophysical properties of galaxies in the early Universe, and their impact on high-redshift observables. We use multi-wavelength observational data including the global 21-cm signal and power spectrum limits from SARAS 3 and HERA respectively, present-day diffuse X-ray and radio backgrounds, and UV luminosity functions (UVLFs) from HST and JWST in the range $z=6\,$ 14.514.5\,14.514.5 to derive our constraints. We constrain a flexible model of halo-mass and redshift dependent star-formation efficiency (SFE), defined as the gas fraction converted into stars, and find that it is best described by little to no redshift evolution at $z\approx 6\,$ 1010\,1010 and rapid evolution at $z\approx 10\,$ 1515\,1515. We derive Bayesian functional posterior distributions for the SFE across this redshift range, inferring that a halo of mass $M_{h}=10^{10}\,\mathrm{M}_{\odot}$ has an efficiency of 222\,2 $\,3\%$ at $z\lesssim 10$, $12\%$ at $z=12$ and $26\%$ at $z=15$. We also find, through synergy between SARAS 3 and UVLFs, that the minimum circular velocity for star-formation in halos is $V_{c}=16.9^{+25.7}_{-9.5}\,\mathrm{km}\,\mathrm{s}^{-1}$ or equivalently $\log_{10}(M_{\text{crit}}/\,\mathrm{M}_{\odot})=8.29^{+1.21}_{-1.08}$ at $z=6$. Alongside these star-formation constraints, we find the X-ray and radio efficiencies of early galaxies to be $f_{X}=0.5^{+6.3}_{-0.3}$ and $f_{r}\lesssim 11.7$ respectively, improving upon existing works that do not use UVLF data. Our results demonstrate the critical role of UVLFs in constraining the early Universe, and its synergies with 21-cm observations, alongside other multi-wavelength observational datasets.
@@ 1 Introduction
The early Universe remains an exciting frontier due to its observational and theoretical prospects. Observations of the Cosmic Microwave Background (CMB) at redshift $z\approx 1100$ ($t_{\text{age}}\sim 0.4\,\mathrm{Myr}$ after the Big Bang) have provided us with a snapshot of the Universe in its infancy, while large-scale spectroscopic surveys like BOSS (Dawson et al., 2013) and DESI (DESI Collaboration et al., 2024) are mapping the nearby Universe in unprecedented detail. The epochs intervening these bookends of our cosmic history, however, remain largely unexplored. Recently, the James Webb Space Telescope (JWST; Gardner et al., 2006) has opened a new window into the distant Universe, with its unprecedented sensitivity and resolution, allowing us to probe some of the earliest galaxies at $z\gtrsim 10$ ($t_{\text{age}}\lesssim 500\,\mathrm{Myr}$) deep into the Epoch of Reionization (EoR). Prior to the JWST, the Hubble Space Telescope (HST) had provided us with a wealth of information on the high-redshift Universe, including the UV luminosity functions (UVLF) of galaxies and the cosmic star-formation history (e.g. Bouwens et al., 2021a).
@@
Now, in the era of the JWST, hundreds of candidate galaxies have been observed at very early epochs (e.g. Castellano et al., 2023; Austin et al., 2023; Leung et al., 2023; Harikane et al., 2023a; Hainline et al., 2024; Finkelstein et al., 2024; Whitler et al., 2025; Pérez-González et al., 2025), with the highest spectroscopically confirmed galaxy JADES-GS-z14-0 at $z=14.32$ (Carniani et al., 2024), alongside a host of distant supermassive black holes, bright AGN and quasars (Larson et al., 2023; Furtak et al., 2024; Yue et al., 2024; Bogdán et al., 2024; Maiolino et al., 2024). Indeed, the abundance of these bright galaxies at such high redshifts given the small survey volumes exceed those expected by simple extrapolations of the UVLF (e.g. Figure 10 of Finkelstein et al. 2024; or Figure 6 of Whitler et al. 2025), or based on theoretical predictions (e.g. Figure 17 of Harikane et al. 2023a; or Figure 9 of Whitler et al. 2025). Several suggestions have been made in literature to resolve this discrepancy, including a top-heavy initial mass function (IMF) of early stars (e.g. Inayoshi et al., 2022; Wang et al., 2023; Trinca et al., 2024; Hutter et al., 2024), an increased star-formation efficiency via bursts (Dekel et al., 2023; Li et al., 2024), the presence of hidden AGN (Hegde et al., 2024), or star-formation variability/stochasticity (e.g. Mason et al., 2023; Mirocha & Furlanetto, 2023; Sun et al., 2023; Pallottini & Ferrara, 2023; Shen et al., 2023).
@@
Despite its grand successes, the era of the very first generation of stars at $z\sim 20\,$ 5050\,5050 ($t_{\text{age}}\sim 50\,$ $\,200\,\mathrm{Myr}$), the so-called metal-free Population III stars (Pop III), will likely be too faint for JWST to detect. The cosmological 21-cm signal, on the other hand, provides a powerful probe of this epoch. The signal is a unique signature of neutral hydrogen in the intergalactic medium (IGM) that is sensitive to the thermal and ionization history of the Universe (Furlanetto et al., 2006; Pritchard & Loeb, 2012) resulting from an interplay between the radiation sourced by stars, stellar remnants, or exotic sources. In particular, the 21-cm signal is sensitive to the properties of the CMB/radio background, early Lyman-$\alpha$ sources such as Pop III stars at Cosmic Dawn ($z\lesssim 30$, e.g. Yajima & Khochfar, 2015; Schauer et al., 2019; Mebane et al., 2020; Gessey-Jones et al., 2022, 2025), heating of the IGM through X-ray photons and Lyman-alpha scattering ($z\lesssim 20$, e.g. Chen & Miralda-Escudé, 2004; Pritchard & Furlanetto, 2007; Fialkov et al., 2014; Pacucci et al., 2014; Reis et al., 2021), and reionization from stellar/quasar ultraviolet emissions ($z\lesssim 10$, e.g. Madau et al., 1997; Ciardi & Madau, 2003)
@@
Recently, Pochinda et al. (2024) conducted the most complete work to date utilizing multi-wavelength synergies alongside the 21-cm signal to constrain properties of the early Universe. Their work uses the SARAS 3 non-detection of the 21-cm global signal (Singh et al., 2022), the latest HERA upper limits on the 21-cm power spectrum (HERA Collaboration et al., 2023), and collated measurements of the cosmic radio background (CRB; Dowell & Taylor, 2018) and cosmic X-ray background (CXB; Hickox & Markevitch, 2006; Harrison et al., 2016). In this work, we extend their analysis by exploring a more flexible star-formation model, and include the most comprehensive UVLF measurements from HST and JWST available to date as an additional observable. In particular, we constrain the star-formation efficiency of galaxies (SFE), an important quantity widely used in analytic galaxy formation models (e.g. Sun & Furlanetto, 2016; Tacchella et al., 2018; Sipple & Lidz, 2024), as well as other astrophysical parameters. In Section 2, we discuss the theoretical framework, including the enhanced SFE model introduced in ths work, followed by an introduction of the observational data used in Section 3. We then describe our methodology for Bayesian inference in Section 4, and present our results in Section 5. We conclude in Section 6.
@@ 2 Theory of the early Universe
In order to simulate the observables from the early Universe, we use the code 21-cm Semi-numerical Predictions Across Cosmic Epochs (21cmSPACE) [^2] For an overview, see cosmicdawnlab.com/21cmSPACE. (Fialkov et al., 2012; Visbal et al., 2012). We briefly describe the simulations in Section 2.1, but otherwise refer the readers to the most recent code development papers for details (Gessey-Jones et al., 2023, 2025). In this work, we add two new features to the code: an upgraded Pop II star-formation efficiency model (described in Section 2.2), and a calculation of UVLFs (described in Section 2.4) for constraining astrophysics using HST/JWST data.
@@ 2.1 Overview of 21cmSPACE
The simulations begin at $z_{\text{max}}=50$ with cosmological initial conditions generated using CAMB (Lewis & Challinor, 2011) for large-scale matter overdensity $\delta$ and baryon-dark matter relative velocity/streaming velocity $v_{\text{bc}}$, and Recfast (Seager et al., 2011) for gas kinetic temperature $T_{\mathrm{K}}$ and ionized fraction $x_{\text{e}}$ assuming Planck 2013 best-fit $\Lambda$CDM cosmology (Planck Collaboration et al., 2014). The initial conditions are set over a cosmological box split into $128^{3}$ cubic cells, assuming periodic boundary conditions, and a resolution of $3$ comoving Mpc, giving a total volume of $(384\,\mathrm{cMpc})^{3}$. The fields $\delta$ and $v_{\text{bc}}$ are evolved forward in time using linear perturbation theory (see, e.g. Barkana, 2016), while $T_{K}$ and $x_{\text{e}}$ are evaluated at each step as we describe below. Halo formation is modelled analytically in each pixel using a hybrid dark matter (DM) halo mass function $dn(M_{h},z)/dM_{h}$ (HMF; Barkana & Loeb, 2004; Tseliakhovich et al., 2011; Fialkov et al., 2012) that depends on the large-scale density fluctuations and streaming velocity resolved at $3\,\mathrm{cMpc}$ scales. This overdensity is capped at $\delta_{\text{max}}=1.3$, much below the critical $\delta_{\text{crit}}\approx 1.69$ to ensure the HMF’s validity. [^3] In order to derive this bound, we first calculate the conditional HMF $\psi(M_{h},z)=dn(M_{h},z|\delta,v_{\text{bc}}=0)/dM_{h}$ for $\delta\in\left[-1,2\right]$, and numerically find the values of $\delta_{\text{max}}$ at which the $d\psi/dM_{h}>0$ for any given $M_{h}$ (i.e., the HMF increases, instead of decreasing, with increasing halo mass which is unphysical behavior). A value of $\delta_{\text{max}}=1.3$ ensures physical validity for all redshifts $z\geq 6$. In the simulation, no pixels reach this threshold at $z=15$, $5$ pixels (0.0002%) at $z=10$, and 3387 pixels (0.16%) at $z=6$, which is a reasonable approximation expected in a semi-numerical model.
@@
We model star-formation to begin with the formation of Pop III stars in metal-free mini-halos in a single burst (per halo) when they reach the critical mass for star-formation $M_{\text{crit}}$ (described in more detail in Section 2.2), following the semi-analytic prescription of Magg et al. (2022). After the first generation of stars, halos undergo a period of recovery $t_{\text{delay}}$ due to Pop III supernovae, as the primordial gas is metal-enriched and the gas re-collapses. The metal-enriched halos are then assumed to form Pop II stars continuously with a star-formation rate density (SFRD) modelled using the sub-grid prescription:
[…]
@@
where $\mathbf{x}$ encapsulates the spatial variation of $\delta$, $v_{\text{bc}}$ and $M_{\text{crit}}$. Here, $t_{\star,\mathrm{II}}=0.2$ is the fraction of Hubble time over which Pop II stars are assumed to form (corresponding to the typical dynamical time for DM halos, Park et al., 2019; Reis et al., 2022) [^4] The dynamical (half-crossing) time for DM halo is ${t_{\text{DM,dyn}}=V_{c}/R_{\text{vir}}=\sqrt{3/(4\pi G\rho_{\text{vir}})}=\sqrt{6/(\Delta_{c}H(z)^{2})}}$ for $\rho_{\text{vir}}=\Delta_{c}\rho_{\text{crit}}$. Different halo radius/mass definitions depending on the overdensity threshold ${\Delta_{c}=50-500}$ (e.g. White, 2001) give values ${t_{\text{DM,dyn}}H(z)\sim 0.10-0.35}$. The Schmidt law (Schmidt, 1959), extended to general star-formation in proto-galaxies (Silk, 1997), yields the star-formation (or gas depletion) timescale ${t_{\text{SF,dyn}}=f_{\star}\rho_{\text{gas}}/\dot{\rho}_{\text{star}}}$. The two are closely linked via gas accretion rate onto galaxies in halos so that $t_{\text{DM,dyn}}\sim t_{\text{SF,dyn}}$ (see, e.g., analytic models of Davé et al., 2012; Dekel & Mandelker, 2014, which are in agreement with simulations)., $f_{\mathrm{II}}$ is mass fraction of halos forming Pop II stars which depends on $t_{\text{delay}}$ (Magg et al., 2022), $f_{\text{g}}/f_{\text{b}}$ is the gas-to-baryon mass fraction affected by $v_{\text{bc}}$ in small halos (Gnedin, 2000; Naoz et al., 2009; Tseliakhovich et al., 2011), and $f_{\star,\mathrm{II}}$ is the Pop II star-formation efficiency (the quantity of interest in this work). Since we intend to focus on Pop II stars, Pop III star-formation parameters are fixed to $f_{\star,\mathrm{III}}=0.002$, with a log-flat IMF in the mass range $2\,\mathrm{M}_{\odot}$ to $180\,\mathrm{M}_{\odot}$ (Gessey-Jones et al., 2022), and a $t_{\text{delay}}=30\,\mathrm{Myr}$ recovery time (e.g. Chiaki et al., 2018; Magg et al., 2022).
@@
Star-formation in halos is modelled through the analytic means described above. However, the radiative fields from stellar and galactic emission propagate to cosmological distances larger than the pixel size and thus need to be handled numerically. This is done by calculating the emissivity $\epsilon(\mathbf{x},\nu)$ of the various radiative species, and propagating them through window functions, taking into account the lightcone effect and redshifting of source spectra. The sources are modelled as follows:
@@
  • X-rays: Emitted by Pop III and Pop II XRBs using the SEDs from Fragos et al. (2013a), as we shall describe in Section 2.3.2. We do not differentiate between the two X-ray populations here, but this distinction will be made in a future work (e.g. Gessey-Jones et al., 2025).
@@
  • Lyman: Lyman-$\alpha$ and Lyman-Werner bands from Pop II stars as described in Barkana & Loeb (2005) calculated using Starburst99 (Leitherer et al., 1999) and from Pop III stars using the spectra from Gessey-Jones et al. (2022), accounting for Ly-$\alpha$ heating and multiple scattering effects (Reis et al., 2021).
@@
  • Radio: Emitted by Pop II and Pop III galaxies using the power-law spectrum from Gürkan et al. (2018); Mirocha & Furlanetto (2019), which we will describe in Section 2.3.3. Here again, we do not differentiate radio emission between the two populations in this work.
@@
    where $\zeta_{\text{ion}}$ is the ionization rate per baryon, $R_{\text{mfp}}=50\,\mathrm{cMpc}$ is the maximum mean free path of ionizing photons (based on end of EoR QSO observations; Wyithe & Loeb, 2004; Zhu et al., 2023), $f_{\text{coll}}(\mathbf{x},R)$ is the collapse fraction of baryons into galaxies averaged over a volume of radius $R$ centered at $\mathbf{x}$, and $x_{\text{e,oth}}(\mathbf{x},R)$ is the fractional ionization from long-range agents like X-rays. If the criterion in Equation 2 is not satisfied down to cell resolution (i.e. $3\,\mathrm{cMpc}$), the cell is treated as a two-phase medium with an ionized and non-ionized fraction (in agreement with cosmological radiative transfer algorithms, Zahn et al., 2011).
@@
Figure 1: Schematic of the Pop II star-formation efficiency model $f_{\star,\mathrm{II}}(M_{h},z)$ used in this work (purple-brown gradient), described in Equation 7, as a function of halo mass $M_{h}$ and redshift $z$. The model is a four-parameter function that is zero below the critical mass $M_{\text{crit}}$ (set to $10^{7}\,\mathrm{M}_{\odot}$ here for illustration, but can be as low as $10^{5}\,\mathrm{M}_{\odot}$ at $z=30$), logarithmically increasing in the molecular cooling regime (light orange shaded region), constant in the atomic cooling regime (dark orange shaded region), and a power-law in the high mass regime (purple region). The turning point ${M_{\text{high}}=M_{0}\left[(1+z)/7\right]^{-\beta_{\star}}}$ is redshift dependent and the evolution rate $\beta_{\star}$ captures the enhanced star-formation efficiency at fixed halo mass in the early Universe. The SFE is naturally capped at $1$. Previous works using 21cmSPACE (e.g. Pochinda et al., 2024) utilize the log cutoff model (dashed orange line), where the SFE is a fixed, constant value in the atomic cooling regime which is too simplistic for the observed mass-dependency of SFE for high mass halos. The single power law (dashed-dotted green line) on the other hand fits provides a better fit to large halos, but assumes a low SFE in unobserved faint, low mass halos which may not be accurate. Finally, the double power law (solid pink line) accounts for feedback from AGN in the highest mass halos, which are abundant at $z\leq 6$, but their impact is less clear at high redshifts.

### v2

@@ Abstract
In the current era of JWST, we continue to uncover a wealth of information about the Universe deep into the Epoch of Reionization. In this work, we use a suite of simulations with 21cmSPACE  to explore the astrophysical properties of early galaxies and their imprint on high-redshift observables. Our analysis incorporates a range of multi-wavelength datasets including UV luminosity functions (UVLFs) from HST and JWST spanning z=6z=6\, 14.5\,14.5, the 21-cm global signal and power spectrum limits from SARAS 3 and HERA respectively, as well as present-day diffuse X-ray and radio backgrounds. We constrain a flexible halo-mass and redshift dependent model of star-formation efficiency (SFE), defined as the fraction of gas converted into stars, and find that it is best described by minimal redshift evolution at $z\approx 6\,$ 10\,10, followed by rapid evolution at $z\approx 10\,$ 15\,15. Using Bayesian inference, we derive functional posteriors of the SFE, inferring that halos of mass $M_{h}=10^{10}\,\mathrm{M}_{\odot}$ have efficiencies of $\approx 1\,$ $2\%$ at $z\lesssim 10$, $\approx 8\%$ at $z=12$ and $\approx 21\%$ at $z=15$. We also highlight the synergy between UVLFs and global 21-cm signal from SARAS 3 in constraining the minimum virial conditions required for star-formation in halos. In parallel, we find the X-ray and radio efficiencies of early galaxies to be $f_{X}=0.8^{+9.7}_{-0.4}$ and $f_{r}\lesssim 16.9$ respectively, improving upon previous works that exclude UVLF data. Our results underscore the critical role of UVLFs in constraining early galaxy properties, and their synergy with 21-cm and other multi-wavelength observations.
@@ 1 Introduction
The early Universe remains an exciting frontier due to its observational and theoretical prospects. Observations of the Cosmic Microwave Background (CMB) at redshift $z\approx 1100$ ($t_{\text{age}}\sim 0.4\,\mathrm{Myr}$ after the Big Bang) have provided us with a snapshot of the Universe in its infancy, while large-scale spectroscopic surveys like Baryon Oscillation Spectroscopic Survey (BOSS; Dawson et al., 2013) and Dark Energy Spectroscopic Instrument survey (DESI; DESI Collaboration et al., 2024) are mapping the nearby Universe in unprecedented detail. The epochs intervening these bookends of our cosmic history, however, remain largely unexplored. Recently, the James Webb Space Telescope (JWST; Gardner et al., 2006) has opened a new window into the distant Universe, with its high sensitivity and resolution, allowing us to probe some of the earliest galaxies at $z\gtrsim 10$ ($t_{\text{age}}\lesssim 500\,\mathrm{Myr}$) deep into the Epoch of Reionization (EoR; e.g. Whitler et al., 2025). Prior to the JWST, the Hubble Space Telescope (HST) had provided us with a wealth of information on the high-redshift Universe, including the UV luminosity functions (UVLF) of galaxies and the cosmic star-formation history (e.g. Bouwens et al., 2021a).
@@
Now, in the era of the JWST, hundreds of candidate galaxies have been observed at very early epochs (e.g. Castellano et al., 2023; Austin et al., 2023; Leung et al., 2023; Harikane et al., 2023a; Hainline et al., 2024; Finkelstein et al., 2024; Whitler et al., 2025; Pérez-González et al., 2025), with the highest spectroscopically confirmed galaxy JADES-GS-z14-0 at $z=14.32$ (Carniani et al., 2024), alongside a host of distant supermassive black holes, bright AGN and quasars (Larson et al., 2023; Furtak et al., 2024; Yue et al., 2024; Bogdán et al., 2024; Maiolino et al., 2024). Indeed, the abundance of these bright galaxies at such high redshifts given the small survey volumes exceed those expected by simple extrapolations of the UVLF (e.g. Figure 10 of Finkelstein et al. 2024; or Figure 6 of Whitler et al. 2025), or based on theoretical predictions (e.g. Figure 17 of Harikane et al. 2023a; or Figure 9 of Whitler et al. 2025). Several suggestions have been made in the literature to resolve this discrepancy, including a top-heavy initial mass function (IMF) of early stars (e.g. Inayoshi et al., 2022; Wang et al., 2023; Trinca et al., 2024; Hutter et al., 2025; Mauerhofer et al., 2025), an increased star-formation efficiency (Dekel et al., 2023; Li et al., 2024; Mauerhofer et al., 2025; Boylan-Kolchin, 2025; Somerville et al., 2025), the presence of hidden AGN (Hegde et al., 2024), or star-formation variability/stochasticity (e.g. Mason et al., 2023; Mirocha & Furlanetto, 2023; Sun et al., 2023a, b; Shen et al., 2023; Kravtsov & Belokurov, 2024; Gelli et al., 2024).
@@
Despite its grand successes, the era of the very first generation of stars at $z\sim 20\,$ 50\,50 ($t_{\text{age}}\sim 50\,$ $200\,\mathrm{Myr}$), the so-called metal-free Population III stars (Pop III), will likely be too faint for JWST to detect. The cosmological 21-cm signal, on the other hand, is a powerful probe of this epoch. The signal is a unique signature of neutral hydrogen in the intergalactic medium (IGM) that is sensitive to the thermal and ionization history of the Universe (Furlanetto et al., 2006; Pritchard & Loeb, 2012) resulting from an interplay between the radiation sourced by stars, stellar remnants, or exotic sources. In particular, the 21-cm signal is sensitive to the properties of the CMB/radio background, early Lyman-$\alpha$ sources such as Pop III stars at Cosmic Dawn ($z\lesssim 30$, e.g. Yajima & Khochfar, 2015; Schauer et al., 2019; Mebane et al., 2020; Gessey-Jones et al., 2022, 2025), heating of the IGM through X-ray photons and Lyman-alpha scattering ($z\lesssim 20$, e.g. Chen & Miralda-Escudé, 2004; Pritchard & Furlanetto, 2007; Fialkov et al., 2014; Pacucci et al., 2014; Reis et al., 2021), and reionization from stellar/quasar ultraviolet emissions ($z\lesssim 10$, e.g. Madau et al., 1997; Ciardi & Madau, 2003)
@@
Recently, Pochinda et al. (2024) conducted the most complete work to date utilizing multi-wavelength synergies alongside the 21-cm signal to constrain properties of the early Universe. Their work uses the SARAS 3 non-detection of the 21-cm global signal (Singh et al., 2022), the latest HERA upper limits on the 21-cm power spectrum (HERA Collaboration et al., 2023), and collated measurements of the cosmic radio background (CRB; Dowell & Taylor, 2018) and cosmic X-ray background (CXB; Hickox & Markevitch, 2006; Harrison et al., 2016). In this work, we extend their analysis by exploring a more flexible star-formation model, and include the UVLF measurements from large HST and JWST surveys as an additional observable. In particular, we constrain the star-formation efficiency of galaxies (SFE), an important quantity widely used in analytic galaxy formation models (e.g. Sun & Furlanetto, 2016; Tacchella et al., 2018; Sipple & Lidz, 2024), as well as the efficiency of X-ray and radio sources. In Section 2, we discuss the theoretical framework, including the enhanced SFE model introduced in this work, followed by an introduction of the observational data used in Section 3. We then describe our methodology for Bayesian inference in Section 4, and present our results in Section 5. We conclude in Section 6.
@@ 2 Theory of the early Universe
[…]
@@ 2.1 Overview of 21cmSPACE
The simulations begin at $z_{\text{max}}=50$ with cosmological initial conditions generated using CAMB (Lewis & Challinor, 2011) for large-scale Eulerian matter overdensity $\delta$ and baryon-dark matter relative velocity/streaming velocity $v_{\text{bc}}$, and Recfast (Seager et al., 2011) for gas kinetic temperature $T_{\mathrm{K}}$ and ionized fraction $x_{\text{e}}$. The initial conditions are set over a cosmological box split into $128^{3}$ cubic cells, assuming periodic boundary conditions, and a resolution of $3$ comoving Mpc, giving a total volume of $(384\,\mathrm{cMpc})^{3}$. The fields $\delta$ and $v_{\text{bc}}$ are evolved forward in time using linear perturbation theory (see, e.g. Barkana, 2016), while $T_{K}$ and $x_{\text{e}}$ are evaluated at each step as we describe below. Halo formation is modelled analytically in each pixel using a hybrid dark matter (DM) halo mass function $dn(M_{h},z|\delta,v_{\text{bc}})/dM_{h}$ (HMF; Barkana & Loeb, 2004; Tseliakhovich et al., 2011; Fialkov et al., 2012) that depends on the large-scale density fluctuations and streaming velocity resolved at $3\,\mathrm{cMpc}$ scales. This large-scale overdensity is capped at $\delta_{\text{max}}=1.3$, which is technically quasi-linear but still below the critical overdensity $\delta_{\text{crit}}\approx 1.69$ often used in HMF formalisms. This particular value was chosen to preserve physical validity of the hybrid HMF we employ. [^3] In order to derive this bound, we first calculate the conditional HMF $\psi(M_{h},z)=dn(M_{h},z|\delta,v_{\text{bc}}=0)/dM_{h}$ for $\delta\in\left[-1,2\right]$, and numerically find the values of $\delta_{\text{max}}$ at which the $d\psi/dM_{h}>0$ for any given $M_{h}$ (i.e., the HMF increases, instead of decreasing, with increasing halo mass which is unphysical behavior). A value of $\delta_{\text{max}}=1.3$ ensures physical validity for all redshifts $z\geq 6$. In the simulation, no pixels reach this threshold at $z=15$, 55 pixels (0.0002%) at $z=10$, and 3387 pixels (0.16%) at $z=6$, which is a reasonable approximation expected in a semi-numerical model.
@@
We model star-formation to begin with the formation of Pop III stars in metal-free mini-halos following the semi-analytic prescription of Magg et al. (2022). In each halo, Pop III stars are assumed to form in a single burst when the halo reaches the critical mass for star-formation $M_{\text{crit}}$ (described in more detail in Section 2.2), with a constant, fixed efficiency of $f_{\star,\mathrm{III}}$ and with stellar masses distributed according to the initial mass function (IMF; Gessey-Jones et al., 2022, 2025). After the first generation of stars reach the end of their lives, halos undergo a period of recovery $t_{\text{delay}}$ due to Pop III supernovae during which they do not form stars. Eventually, the metal-enriched gas re-collapses, and the halos start forming Pop II stars. We assume that Pop II stars form continuously with a star-formation rate density (SFRD) modelled using the sub-grid prescription:
$\begin{split}&\dot{\rho}_{\star,\rm II}(\mathbf{x},z)=\big{(}1+\delta(\mathbf{x})\big{)}\,\,\times\,\,\frac{1}{t_{\star,\mathrm{II}}H(z)^{-1}}\,\,\times\\ &\int_{M_{\text{crit}}}^{M_{\text{max}}}\left({f_{\star,\mathrm{II}}}(\mathbf{x})f_{\mathrm{II}}(\mathbf{x},t_{\text{delay}})\frac{f_{\text{g}}}{f_{\text{b}}}(\mathbf{x},M_{h})\dfrac{\Omega_{\text{b}}}{\Omega_{\text{m}}}M_{h}\right)\dfrac{dn(\mathbf{x},M_{h},z)}{dM_{h}}dM_{h}\end{split}$ (1)
@@
where $\mathbf{x}$ is the spatial coordinate in the simulation box (with varying $\delta$, $v_{\text{bc}}$ and $M_{\text{crit}}$), and $(1+\delta)$ accounts for conversion from Lagrangian to Eulerian space (see, e.g. Mesinger et al., 2011; Tseliakhovich et al., 2011; Muñoz, 2023). Here, $t_{\star,\mathrm{II}}=0.2$ is the fraction of Hubble time over which Pop II stars are assumed to form (corresponding to the typical dynamical time for DM halos, Park et al., 2019; Reis et al., 2022b) [^4] The dynamical (half-crossing) time for DM halo is $t_{\text{DM,dyn}}=V_{c}/R_{\text{vir}}=\sqrt{3/(4\pi G\rho_{\text{vir}})}=\sqrt{6/(\Delta_{c}H(z)^{2})}$} for $\rho_{\text{vir}}=\Delta_{c}\rho_{\text{crit}}$. Different halo radius/mass definitions depending on the overdensity threshold $\Delta_{c}=50-500$} (e.g. White, 2001) give values $t_{\text{DM,dyn}}H(z)\sim 0.10-0.35$}. The Schmidt law (Schmidt, 1959), extended to general star-formation in proto-galaxies (Silk, 1997), yields the star-formation (or gas depletion) timescale $t_{\text{SF,dyn}}=f_{\star}\rho_{\text{gas}}/\dot{\rho}_{\text{star}}$}. The two are closely linked via gas accretion rate onto galaxies in halos so that $t_{\text{DM,dyn}}\sim t_{\text{SF,dyn}}$ (see, e.g., analytic models of Davé et al., 2012; Dekel & Mandelker, 2014, which are in agreement with simulations)., $f_{\mathrm{II}}$ is mass fraction of halos forming Pop II stars which depends on $t_{\text{delay}}$ (Magg et al., 2022), $f_{\text{g}}/f_{\text{b}}$ is the gas-to-baryon mass fraction affected by $v_{\text{bc}}$ in small halos (Gnedin, 2000; Naoz et al., 2009; Tseliakhovich et al., 2011), and $f_{\star,\mathrm{II}}$ is the Pop II star-formation efficiency (the quantity of interest in this work). Since we intend to focus on Pop II stars, Pop III star-formation parameters are fixed to $f_{\star,\mathrm{III}}=0.002$, with a log-flat IMF in the mass range $2\,\mathrm{M}_{\odot}$ to $180\,\mathrm{M}_{\odot}$ (Gessey-Jones et al., 2022), and a $t_{\text{delay}}=30\,\mathrm{Myr}$ recovery time (e.g. Chiaki et al., 2018; Magg et al., 2022).
@@
Star-formation in halos is modelled through the analytic means described above. However, the radiative fields from stellar and galactic emission propagate to cosmological distances larger than the pixel size and thus need to be handled numerically. This is done by calculating the comoving emissivity $\epsilon(\mathbf{x},\nu,z)$ of the various radiative species, and propagating them through window functions taking into account the lightcone effect and redshifting of source spectra. The sources are modelled as follows:
@@
  • X-rays: Emitted by Pop III and Pop II high-mass X-ray binaries using the SEDs from Fragos et al. (2013b), as we shall describe in Section 2.3.2. We do not differentiate between the two X-ray populations here, but this distinction will be made in a future work (e.g., following the prescription developed in Gessey-Jones et al., 2025).
@@
  • Lyman: Lyman-$\alpha$ and Lyman-Werner bands from Pop II stars as described in Barkana & Loeb (2005) calculated using Starburst99 (Leitherer et al., 1999) and from Pop III stars using the spectra from Gessey-Jones et al. (2022). We account for Ly-$\alpha$ heating and multiple scattering effects (Reis et al., 2021).
@@
  • Radio: Emitted by Pop II and Pop III galaxies using the power-law spectrum from Gürkan et al. (2018); Mirocha & Furlanetto (2019), which we will describe in Section 2.3.3. Here again, we do not differentiate radio emission between the two populations. We also exclude the effect of CMB heating (Venumadhav et al., 2018) in this work.
@@
    where $\zeta_{\text{ion}}$ is the ionization rate per baryon, $R_{\text{mfp}}=50\,\mathrm{cMpc}$ is the maximum mean free path of ionizing photons (based on end of EoR QSO observations; Wyithe & Loeb, 2004; Qin et al., 2021; Zhu et al., 2023), $f_{\text{coll}}(\mathbf{x},R)$ is the collapse fraction of baryons into galaxies averaged over a volume of radius $R$ centered at $\mathbf{x}$, and $x_{\text{e,oth}}(\mathbf{x},R)$ is the fractional ionization from long-range agents like X-rays. If the criterion in Equation 2 is not satisfied down to cell resolution (i.e. $3\,\mathrm{cMpc}$), the cell is treated as a two-phase medium with an ionized and non-ionized fraction (in agreement with cosmological radiative transfer algorithms, Zahn et al., 2011; Hutter, 2018).
@@
Figure 1: Schematic of the Pop II star-formation efficiency model $f_{\star,\mathrm{II}}(M_{h},z)$ used in this work (purple-brown gradient, at various redshifts), described in Equation 7, as a function of halo mass $M_{h}$ and redshift $z$. The model is a four-parameter function that is zero below the critical mass $M_{\text{crit}}$ (set to $10^{7}\,\mathrm{M}_{\odot}$ here for illustration, but can be as low as $10^{5}\,\mathrm{M}_{\odot}$ at $z=30$), logarithmically increasing in the molecular cooling regime (left, light orange shaded region), constant in the atomic cooling regime (center, dark orange shaded region), and a power-law in the high mass regime (right, purple region). The turning point $M_{\text{high}}=M_{0}\left[(1+z)/7\right]^{-\beta_{\star}}$} is redshift dependent and the evolution rate $\beta_{\star}$ captures the enhanced star-formation efficiency at fixed halo mass in the early Universe. The SFE is naturally capped at $1$. The example illustrated here corresponds to $\tilde{f}_{\star,\mathrm{II}}=0.05,M_{0}=10^{10}\,\mathrm{M}_{\odot},\alpha_{\star}=0.5,\beta_{\star}=4$. Previous works using 21cmSPACE (e.g. Pochinda et al., 2024) utilize the log cutoff model (dashed orange line, shown at $z=6$), where the SFE is a fixed, constant value in the atomic cooling regime which is too simplistic for the observed mass-dependency of SFE for high mass halos. The single power law (dashed-dotted green line) on the other hand provides a better fit to large halos, but assumes a low SFE in unobserved faint, low mass halos which may not be accurate. Finally, the double power law (solid pink line) accounts for feedback from AGN in the highest mass halos, which are abundant at $z\leq 6$, but their impact is less clear at high redshifts.